OpenStudy (anonymous):

Which of the following points is a solution to the system of linear inequalities?

OpenStudy (anonymous):

A. (0, –6) B. (1, –9) C. (2, –11) D. (3, –15)

OpenStudy (chrisplusian):

Do you know how to graph those lines? Pretend they are not inequalities and first graph them.

OpenStudy (chrisplusian):

Are you able to do that part?

OpenStudy (anonymous):

yes, i can do that.

OpenStudy (chrisplusian):

Ok graph one line at a time. Now select a test point

OpenStudy (anonymous):

I don't know or understand how to re-write the equations to graph them.

OpenStudy (chrisplusian):

ok what you have is\[-2x+y <1\]

OpenStudy (chrisplusian):

Replace the inequality with an equal sign. so:\[-2x+y=-1\]

OpenStudy (chrisplusian):

now get that in slope intercept form.

OpenStudy (chrisplusian):

Which would be y=2x-1

OpenStudy (chrisplusian):

can you graph that one now?

OpenStudy (anonymous):

How do you re-write it into slope intercept form? That's my question.

OpenStudy (chrisplusian):

ok You understand what I did right? I replaced the inequality with an equal. Do you see that?

OpenStudy (anonymous):

Yes, I mean once you have -2x+y=1

OpenStudy (chrisplusian):

Ok just checking now...

OpenStudy (chrisplusian):

You add 2x to both sides of the equal sign

OpenStudy (chrisplusian):

(2x)-2x+y=-1+(2x)

OpenStudy (chrisplusian):

when you simplify that you get y=-1+2x and if you rearrange the terms you get y=2x-1

OpenStudy (chrisplusian):

Make sense?

OpenStudy (anonymous):

Okay, I see.

OpenStudy (chrisplusian):

Do you know what to do from here? Or you want me to continue?

OpenStudy (anonymous):

I know how to graph them, and I now understand. If I need more help I'll let you know. Thank you for your help. :)

OpenStudy (chrisplusian):

No problem. Last thing is you can check the point on the graph to see if it satisfies both inequalities. There is a shortcut method to this, but I like to explain where it comes from. The shortcut method is to plug the proposed solutions (from the multiple choice answers) into the inequality and simplify. If you get 6>2 you know that proposed solution did not work. If you can check a point in both inequalities and it gives true statements for both then you have found your match. I hope this has helped. Good luck